OpenStudy (anonymous):

What mass of O2, in grams, is required to completely react with 0.025 g C3H8?

OpenStudy (anonymous):

MEDALS WILL BE REWARDED! The molar mass of oxygen gas (O2) is 32.00 g/mol. The molar mass of C3H8 is 44.1 g/mol. What mass of O2, in grams, is required to completely react with 0.025 g C3H8? 0.018 0.034 0.045 0.091

OpenStudy (abmon98):

C3H8+5O2-->3CO2+4H2O Number of Moles=Mass(g)/Molar Mass(g/mol) 0.025/(12*3)+(8*1) =0.00056818181 moles of C3H8 Look at the mole ratio of C3H8 to O2 1:5 0.00056818181:x x moles is obtained by multiplying by 5 and dividing by 1 0.00284090909 moles of oxygen Molar Mass*Number of Moles=Mass 0.00284090909*(16*2)=Mass 0.00284090909*32=Mass

OpenStudy (abmon98):

So o go with the last choice 0.091

OpenStudy (anonymous):

Good Job Abmon98

OpenStudy (abmon98):

Thanks @Katiria12 :)

OpenStudy (anonymous):

you're welcome @Abmon98 ;)